### v2

@@ 3.5 Measurement and Diagnostics
The battery pack shall report its state of charge and the status of the system components to the vehicle controller. In addition, in some cases, such as an overcurrent, the pack should be able to act appropriately. A combination of cells constitutes a module and a combination of modules forms a pack. The module can have a structure similar to 4S3P, which means that three cells are connected in parallel, and four of these parallel connected cells are connected in series to form a 12 V battery module. Since 2010, there has been a significant challenge in the EV industry related to the lack of a common communication standard for battery packs and BMS across different manufacturers. This absence of a universal standard has led to integration difficulties when trying to connect various EV battery systems to diagnostic tools and other vehicle systems. This lack of standardization has particularly impacted the development of diagnostic tools that can interface with multiple battery packs from different manufacturers. However, significant progress has been made, and tools like AUTEL have become indispensable in the diagnosis of a wide range of EVs. AUTEL stands out due to its ability to interface with numerous battery packs, thanks to its extensive range of adapters that accommodate different connectors and plugs. While improvements in communication standards are ongoing, the industry still faces substantial challenges in achieving universal interoperability across all EV battery systems For each battery module, all cell voltages and temperatures usually need to be measured in the BMS. Typically, there is a current sensor for the entire pack that measures the final output current. This means that there is no need to measure all modules separately, or even measure the individual cell currents that are connected in parallel. The entire pack voltage and current should be measured with an accuracy of approximately 1-3% which provides the possibility of calculating the pack power with an accuracy of approximately 3-5%. Having different accuracy values at different operating points is common. For example, for measurement and estimation, the required accuracy should be higher than the accuracy needed for protection. If the battery pack and consequently the energy, is calculated more accurately, it is equivalent to a greater driving range with one charge. Usually, a 5-10% capacity reserve in a battery is considered for EVs. The new trend is to reduce this number, and companies are now trying to reduce this reserve capacity by better estimating of the system. However, extensive testing, data collection over the years, and aging and environmental conditions, including driving behaviors, provide a more accurate estimation of battery energy over its lifetime. Therefore, it is possible to reduce the reserve energy of the pack, which is an important selling point for customers. In addition to the required measurement accuracy, the system should be able to report its operating status, including faults. If a fault occurs, it is recorded in the system. The diagnostics involve knowing the system status, including subsystems, as much as possible. For example, if the main contactor is welded, the system can detect the fault. In addition, there is a large action list to identify what action should be taken in the case of a failure. For diagnostics, one can mention all cell voltages, main contactor and pre-charge contactor status, auxiliary supply status, current sensor status, connector status (if they are connected or disconnected), and some other similar signals. The diagnostics are nearly correlated with the automotive safety integrity level (ASIL) of the subsystem. As a result of the required ASIL level, there could be different measurement solutions from a mix of sensor and software estimation to a level where a redundant sensor is used. This is discussed in the Safety section. For each measurement point, the required ASIL level was part of the requirement. The ASIL level was determined after performing a full system safety analysis.
@@ 5 Environmental Requirements
6 Battery Thermal Management system (BTMS) and Safety
Battery thermal management and safety is an important aspect of EVs, as an emerging technology. One of the major safety concerns is thermal runaway, which can occur due to excessive heat, internal short circuits, overcharging, or mechanical damage to the battery cells. This can lead to battery pack failure, fire, or even explosion, posing a significant hazard to the vehicle, passengers, and the surrounding environment.
To mitigate the risks associated with thermal runaway, Battery Thermal Management Systems (BTMS) are employed. These systems are designed to maintain safe operating temperatures across the battery pack, preventing overheating and minimizing the likelihood of thermal runaway events. Proper temperature regulation is essential for ensuring the performance, longevity, and safety of battery packs. There are various technologies used within BTMS to control the temperature of the battery pack R3_Comprehesive :
@@ 6 Battery Thermal Management system (BTMS) and Safety
  • Active and Passive Cooling: Active cooling systems use pumps and liquid coolants to remove excess heat from the battery cells, ensuring that the cells remain within their safe temperature range. Passive cooling systems, on the other hand, use natural convection, heat pipes, or other methods to dissipate heat without the use of pumps or external power sources. Both active and passive cooling mechanisms play a vital role in preventing overheating of the battery pack. These systems ensure that the battery operates within safe thermal limits during high-demand scenarios, such as fast charging or high-power discharges, and thus help prevent the initiation of thermal runaway R3_thermal ; R3_Comprehesive ; R3_development .
@@
  • Phase-Change Materials (PCMs): PCMs are substances that absorb and store heat as they change phase (from solid to liquid) at a specific temperature. In a BTMS, PCMs are integrated into the battery pack to help absorb excess heat during periods of high charge or discharge rates. This helps maintain a more uniform temperature across the battery pack and prevents localized overheating. As the temperature rises, the PCM absorbs thermal energy and transitions to a liquid state, stabilizing the pack’s temperature. When the temperature falls, the material solidifies, releasing stored heat. This capability can significantly enhance the safety and efficiency of EV battery packs R3_development .
@@
  • Immersion Cooling: Immersion cooling involves submerging the battery pack in a thermally conductive liquid that directly absorbs heat from the cells. This method of thermal management provides more uniform and efficient cooling compared to traditional air cooling systems. Immersion cooling is especially useful for high-power applications, where rapid temperature fluctuations are more likely. By submerging the cells in a liquid coolant, the heat is more effectively transferred from the cells to the coolant, reducing the chance of localized overheating and thermal runaway R3_thermal .
@@
In addition to these technologies, BMS also play a key role in ensuring safety. They continuously monitor the battery’s voltage, temperature, and other critical parameters, triggering corrective actions when abnormal conditions are detected. For example, the BMS can activate the cooling system or limit charging rates to prevent overheating. There have been several reports of EV accidents Crash . There are different safety layers and standard regulations from the vehicle level to the sub-component level. In response to fulfilling standard regulations and achieving a safe system, the ISO 25265 series has been in place, followed by automakers and suppliers ISO26262-1 ; ISO26262-2 ; ISO26262-3 ; ISO26262-4 ; ISO26262-5 ; ISO26262-6 ; ISO26262-7 ; ISO26262-8 ; ISO26262-9 ; ISO26262-10 ; ISO26262-11 ; ISO26262-12 . The topic is briefly reviewed in this section, and it is expected that standards or relevant references will be visited for more information.
@@ 7.1 Cell Technology
Li-ions are the dominant chemical species in automotive battery cells. There are different types of chemistries, such as Lithium Cobalt Oxide (LCO), Lithium Iron Phosphate (LFP), Lithium Nickel Cobalt Aluminum Oxide (NCA), and Lithium Nickel Manganese Cobalt Oxide (NMC), to name but a few. However, there are more details in the chemistry that are not publicly available. Fig. 5 shows LCO, LFP, NCA, and NMC battery types, which are different in the cathode types and have different nominal voltage, and so on. The charactristics of these batteries are presented in Table 2.
@@
NMC and LFP cells are widely used in automotive applications. The LFP has a lower voltage but better temperature stability. Battery cells are available mainly in three packages: pouch, cylindrical, and prismatic, as shown in Fig. 6. The pouch is easier to produce and to cool in mass and is used by most automakers. In contrast, Tesla uses cylinder cells that require a more complicated cooling structure and better safety. As mentioned earlier, cell chemistries are very different and it is difficult to classify them accurately. Examples of main cell manufacturers for automotive applications are LG, CATL, and Panasonic. More companies and new companies, such as Northvolt, will emerge as well. The ideal cell should be sufficiently high to provide the required ampere. For the above-mentioned 100 kWh pack, the ideal single-cell capacity is 250 Ah. However, the available cell in the automotive is approximately 80 Ah for the pouch type. Hence, three cells were be connected in parallel to provide the required $Ah$. Pouch cells with a capacity of approximately 200 Ah are under development, and it is expected to see them in a couple of years.
@@ 8 Battery Management System (BMS)
The BMS in the battery pack has several functions, including monitoring of cells and the temperature at some points, controlling the contactors, monitoring of the pack current, communication towards the vehicle, and additional functions that will be discussed in this section. The idea was to complement the materials provided in the previous sections. BMS play a critical role in ensuring the optimal performance, safety, and longevity of EV battery packs. Among the primary tasks of the BMS are the SOC estimation, which determines the remaining charge in the battery; SOH estimation, which evaluates the overall health of the battery; and temperature monitoring, which ensures that the battery operates within safe temperature limits. These parameters are crucial for maintaining efficient and reliable operation.
8.1 Estimation methods
@@ 8.1.1 State-of-Charge (SoC) Estimation
The SOC estimation is crucial for ensuring that the battery is neither overcharged nor excessively discharged, both of which can shorten its lifespan. The most common methods for SOC estimation include the following:
@@
    Coulomb-counting Method: This classical method, which can also be called the ampere count method, calculates the SoC by integrating the current over time Larijani . Although simple and computationally efficient, it can suffer from errors due to the accumulation of drift over time, requiring periodic recalibration or corrections. In addition, the initial SoC should be known to accurately compute the SoC over time Edorado_2 .
@@
    Model-Based Methods: Recent advancements have introduced more sophisticated methods, such as those based on Kalman filters, which dynamically adjust to changing conditions. One notable study by Locorotondo et al. Locorotondo demonstrated the use of model-adaptive Kalman filters to estimate SoC more accurately by adjusting the model parameters based on real-time data, such as current and voltage measurements. These methods can reduce the error margins seen in simpler techniques and provide more reliable estimates of SoC.
@@
    Spectroscopy Methods: These methods use techniques such as impedance spectroscopy to estimate SoC by analyzing the battery’s internal resistance and other electrochemical properties spect . While these methods can provide high precision, they tend to be more complex and require specialized equipment.
@@ 8.1.2 State of Health (SoH) Estimation
The SoH refers to the ability of the battery to perform at its rated capacity and maintain charge retention over time. Model-based approaches, similar to those used for SoC estimation, are commonly employed for SoH estimation. These models often incorporate aging effects and degradation patterns, providing more accurate forecasts of battery performance over their lifecycles battery_degradation ,Edorado . The temperature is a critical parameter for maintaining battery safety and performance. Excessive temperatures can accelerate aging, whereas low temperatures can reduce efficiency. The BMS typically include temperature sensors located at multiple points in the battery pack to monitor these fluctuations. Advanced methods for temperature management often involve combining sensor data with predictive models to optimize the battery’s thermal management system in real-time. Despite significant progress, several challenges remain in the accurate estimation of SoC, SoH, and temperature. For example, temperature variations across cells and the aging effects of battery components can complicate estimation processes. Recent advancements in machine learning and adaptive algorithms show promise in improving the accuracy of these estimations. By improving the methods for estimating these parameters, modern BMS contribute significantly to the safety, efficiency, and durability of EV batteries, ensuring that they operate within optimal conditions throughout their lifecycle.
@@ 9 Other Topics and New Trends
The specific energy and power are important parameters in the cell and pack levels that can be countered as performance indices. As mentioned in the previous section, the first step is to construct battery modules as independent units and then integrate them into a pack to fulfill all requirements and standard regulations. The general trend is towards more compact and lighter systems with reduced costs. For pouch cells, an energy density of 650 Wh/L and specific energy of 270 Wh/kg can be found in vehicles. For upcoming cells, a 10-20% enhancement will appear in a couple of years. As an approximation, these numbers can be reduced by approximately 40% at the pack level. For example, an automotive battery pack using cells with 270 Wh/kg, can have an energy density of approximately 180 Wh/kg. New cell technologies with improved energy density or specific energy numbers, more integration in packaging, and improved technology for other parts of the system, such as contactors, sensors, and elimination of wires (using wireless), are needed to reduce the weight and volume of the pack. Hence, new materials and chemistry play key roles in meaningful changes in technological advancement. The battery system voltage is another interesting topic. A nominal voltage of 350 V is widely used for passenger cars. However, in some cases, the system voltage was 700-800 V. The power requirement has increased from less than a hundred kW to more than 200 kW and in some cases, it has increased even further. Consequently, it is necessary to increase the system voltage level to reduce the current level in the system. Consequently, several high-voltage developments such as 800 V is under development by the automotive industry, which will increasingly appear in the market Haghbin2 . Another point that should be mentioned is the charging infrastructure.
@@
Another interesting area of research and development in the battery sector involves sodium-based cells, which present an alternative to lithium-ion batteries. Sodium-based batteries utilize sodium ions instead of lithium ions, offering potential advantages such as abundant raw material availability and lower costs, reducing the environmental and ethical concerns associated with lithium and cobalt mining. These sodium-ion batteries are gaining attention for their potential to power EVs in the future, especially as the demand for sustainable and cost-effective solutions increases. Although still in the early stages compared to lithium-ion batteries, sodium-based cells are expected to play a role in the evolution of EV batteries TNO . Another emerging trend is the development of batteries that do not contain rare earth elements. Many current EV battery chemistries, such as those using lithium cobalt oxide, rely heavily on rare earth materials. However, as the industry moves toward more sustainable practices, researchers are focusing on new technologies that minimize or eliminate the use of rare earths. These technologies aim to reduce environmental impact and alleviate concerns about supply chain limitations and geopolitical issues surrounding the extraction of these critical materials. TNO battery technology, which is characterized by high charging and discharging power capabilities, is also gaining attention. This innovation provides rapid charge/discharge cycles, offering the potential for reducing charging times significantly and improving vehicle performance. As the automotive industry continues to focus on reducing charging times and enhancing power delivery, TNO’s battery technology could become a key player in meeting these demands. Additionally, battery swapping systems, as introduced by companies like NIO, offer a unique solution to the issue of long charging times NIO . Instead of waiting for a battery to charge, NIO offers the ability to swap out the battery entirely in a matter of minutes. This system addresses the challenge of limited fast-charging infrastructure and provides a flexible solution for EV owners, particularly in regions with growing EV adoption but underdeveloped charging networks. Increasing the battery voltage may add issues to the available fast-charging infrastructure. Some fast chargers can not charge 800 V batteries. Hence, there are some works under development to change the classical structure such that the vehicle can utilize the low-voltage charger to charge high-voltage packs. In addition to passenger cars, battery packs are required for buses, trucks, and aviation vehicles. Usually, the solutions developed for passenger cars are optimized, and the results are expected to be utilized in other sections. For aviation electrification, some major hindrances require special attention and further development. The classical structure of a battery pack is inappropriate for aviation. It is not possible to disconnect the battery pack in the case of an extreme fault, and new architectures are required. In addition, the environmental conditions are harsher and packaging may be more challenging. Available aviation solutions, such as satellites, are extremely expensive and not affordable. Hence, more work and maturity are required in this sector.
10 Conclusions

### v3

@@ 3.5 Measurement and Diagnostics
The battery pack shall report its state of charge and the status of the system components to the vehicle controller. In addition, in some cases, such as an overcurrent, the pack should be able to act appropriately. A combination of cells constitutes a module and a combination of modules forms a pack. The module can have a structure similar to 4S3P, which means that three cells are connected in parallel, and four of these parallel connected cells are connected in series to form a 12 V battery module. Since 2010, there has been a significant challenge in the EV industry related to the lack of a common communication standard for battery packs and BMS across different manufacturers. The absence of a universal standard has led to difficulties when connecting various EV battery systems to diagnostic tools and other vehicle systems. Also, it effectively limited the progress of diagnostic tools that can connect with battery packs from various manufacturers. Nevertheless, substantial progress has been made, and tools like AUTEL play a key role in diagnosing a wide range of EVs AUTEL . AUTEL distinguishes itself by its capability to connect with a wide variety of battery packs, because of its wide range of adapters which are designed to support different connectors and plugs. Although improvements in communication standards are ongoing, the industry still encounters significant challenges in achieving universal interoperability across the entire EV battery system. For each battery module, all cell voltages and temperatures typically need to be measured in the BMS. Typically, there is a current sensor for the entire pack that measures the final output current. This means that there is no need to measure all modules separately, or even measure the individual cell currents that are connected in parallel. The entire pack voltage and current should be measured with an accuracy of approximately 1-3% which provides the possibility of calculating the pack power with an accuracy of approximately 3-5%. Having different accuracy values at different operating points is common. For example, for measurement and estimation, the required accuracy should be higher than the accuracy needed for protection. If the battery pack and consequently the energy, is calculated more accurately, it is equivalent to a greater driving range with one charge. Usually, a 5-10% capacity reserve in a battery is considered for EVs. The new trend is to reduce this number, and companies are now trying to reduce this reserve capacity by better estimating the system. However, extensive testing, data collection over the years, and aging and environmental conditions, including driving behaviors, provide a more accurate estimation of battery energy over its lifetime. Therefore, it is possible to reduce the reserve energy of the pack, which is an important selling point for customers. In addition to the required measurement accuracy, the system should be able to report its operating status, including faults. If a fault occurs, it is recorded in the system. The diagnostics involve knowing the system status, including subsystems, as much as possible. For example, if the main contactor is welded, the system can detect the fault. In addition, there is a large action list to identify what action should be taken in the case of a failure. For diagnostics, one can mention all cell voltages, main contactor and pre-charge contactor status, auxiliary supply status, current sensor status, connector status (if they are connected or disconnected), and some other similar signals. The diagnostics are nearly correlated with the automotive safety integrity level (ASIL) of the subsystem. As a result of the required ASIL level, there could be different measurement solutions from a mix of sensor and software estimation to a level where a redundant sensor is used. This is discussed in the Safety section. For each measurement point, the required ASIL level was part of the requirement. The ASIL level was determined after performing a full system safety analysis.
@@ 5 Environmental Requirements
6 Safety
Safety is an important feature of EVs, as an emerging technology, which play a pivotal role in the EV’s performance. Thermal runaway is one of the major concerns of safety, which typically occurs due to excessive heat, internal short circuits, overcharging, and mechanical damage to the battery cells. If these things happen, it causes the failure of the battery pack, fire, or even explosion, which poses a significant hazard to the EV and passengers. In this regard, to mitigate the possible problems associated with thermal runaway, the BMS is used, which is designed to maintain safe operating temperatures of the battery pack, preventing overheating and minimizing the probability of thermal runaway events. Proper temperature regulation is a crucial factor in the performance, longevity, and safety of battery packs. There are various technologies used within BMS to control the temperature of the battery pack R3_Comprehesive :
@@ 6 Safety
  • Active and Passive Cooling: Active cooling systems use pumps and liquid coolants to remove excess heat from the battery cells, ensuring that cells operate in the safe temperature range. On the contrary, passive cooling systems use natural convection, heat pipes, or other methods to dissipate heat flow without the use of external power sources. Nevertheless, both active and passive cooling mechanisms are essential in preventing overheating of the battery pack. These systems prevent the starting thermal runaway and ensure that the battery temperature is in the safe thermal limits during high-demand scenarios, like fast charging or discharging with high power R3_thermal ; R3_Comprehesive ; R3_development .
@@
  • Phase-Change Materials (PCMs): PCMs are substances which absorb heat while they change phase from solid to liquid at a specific temperature. Some BMS use the PCM to help absorb excess heat during high charge or discharge C-rates. By doing so, a more uniform temperature across the battery pack can be obtained, which prevents the hot spot due to overheating. When the temperature rises, the PCM absorbs the energy and transitions to a liquid state, thereby stabilizing the battery pack’s temperature. When the temperature falls, the material solidifies, releasing stored heat. This capability can extremely enhance the safety and efficiency of the EV battery packs R3_development .
@@
  • Immersion Cooling: Immersion cooling involves submerging the battery pack inside a thermally conductive liquid that directly absorbs heat from the cells. As compared to air cooling systems, this method provides more uniform and efficient cooling. Immersion cooling is especially useful for high-power applications, in which fast temperature fluctuations are more likely. By submerging the cells in a liquid coolant, the heat is more effectively transferred from the cells to the coolant, reducing the possibility of thermal runaway or having the hot spot R3_thermal .
@@
Besides these technologies, BMS also plays a crucial role in safety by monitoring the battery’s voltage, doing corrective action when abnormal conditions are detected, and other critical parameters. For example, the BMS can enable the cooling system or limit charging rates to prevent overheating if it is required. There have been several reports of EV accidents Crash . There are different safety layers and standard regulations from the vehicle level to the sub-component level. In response to fulfilling standard regulations and achieving a safe system, the ISO 25265 series has been established, followed by automakers and suppliers ISO26262-1 ; ISO26262-2 ; ISO26262-3 ; ISO26262-4 ; ISO26262-5 ; ISO26262-6 ; ISO26262-7 ; ISO26262-8 ; ISO26262-9 ; ISO26262-10 ; ISO26262-11 ; ISO26262-12 . The topic is briefly reviewed in this section, and it is expected that standards or relevant references will be visited for more information.
@@ 7.1 Cell Technology
Li-ions are the dominant chemical species in the automotive battery cells. There are different types of common li-ions batteries, such as Lithium Cobalt Oxide (LCO), Lithium Iron Phosphate (LFP), Lithium Nickel Cobalt Aluminum Oxide (NCA), and Lithium Nickel Manganese Cobalt Oxide (NMC), to name but a few. Fig. 7 shows the LCO, LFP, NCA, and NMC battery types, which are different in the cathode types and have different nominal voltage, mass, energy density, and so on. The characteristics of these batteries are presented in Table 2.
@@
NMC and LFP cells are widely used in automotive applications. The LFP has a lower voltage but better temperature stability Americans . Battery cells are available mainly in three packages: pouch, cylindrical, and prismatic, as shown in Fig. 8. The pouch is easier to produce and to cool in mass and is used by most automakers. In contrast, Tesla uses cylinder cells that require a more complicated cooling structure and better safety. As mentioned earlier, cell chemistries are very different, and it is difficult to classify them accurately. Examples of main cell manufacturers for automotive applications are LG, CATL, and Panasonic. More companies and new companies, such as Northvolt, will emerge as well. The ideal cell should be sufficiently high to provide the required ampere. For the above-mentioned 100 kWh pack, the ideal single-cell capacity is 250 Ah. However, the available cell in the automotive industry is approximately 80 Ah for the pouch type. Hence, three cells were be connected in parallel to provide the required $Ah$. Pouch cells with a capacity of approximately 200 Ah are under development, and it is expected to see them in a couple of years.
@@ 8 Battery Management System (BMS)
The BMS in the battery pack has several functions, including monitoring of cells and the temperature at some points, controlling the contactors, monitoring of the pack current, communication towards the vehicle, and additional functions that will be discussed in this section. The idea was to complement the materials provided in the previous sections. BMS plays a critical role in ensuring optimal performance, safety, and longevity of EV battery packs. Among the primary tasks of the BMS are SOC estimation, which determines the remaining charge in the battery; SOH estimation, which evaluates the overall health of the battery; and temperature monitoring, which ensures that the battery operates within safe temperature limits. These parameters are crucial for maintaining efficient and reliable operation.
8.1 Battery SoC and SoH estimations
@@ 8.1.1 State-of-Charge (SoC) Estimation
The SOC estimation determines the remaining charge of the battery cell, which plays a key role to ensure that the battery is neither overcharged nor excessively discharged, both of which can shorten the battery cell life-time. The most common methods for SOC estimation include the following:
@@
    Coulomb-counting Method: This method, can also be called the ampere count method, calculates the battery cell SoC by integrating its current over time Xiong ; Larijani . Although this method is simple, it can suffer from errors of the accumulation of drift over time, requiring periodic corrections. Furthermore, the initial SoC should be known to accurately compute the SoC over time Edorado_2 .
@@
    Model-Based Methods: This method uses the battery cell model, for which there are recent advancements too, such as those method which are based on Kalman filters, which dynamically adjust to variable conditions. The study by Locorotondo et al. Locorotondo presented the use of model-adaptive Kalman filters to estimate SoC more accurately by adjusting the model parameters based on real-time data, like current and voltage measurements. These methods can reduce the error margins which has been seen in simpler techniques, thereby providing more reliable estimates of SoC.
@@
    Spectroscopy Methods: These methods use techniques like impedance spectroscopy to estimate SoC by analyzing the internal resistance of the battery and other electrochemical properties spect . Although these methods have high precision, actually they are a little complex and require specialized equipment, and are not suitable for real-time implementation.
@@ 8.1.2 State of Health (SoH) Estimation
SoH evaluates the overall health of a battery which refers to the ability of the battery to perform at the rated capacity and maintain charge retention over time. Model-based approaches, which are similar to those used for SoC estimation, are commonly employed for SoH estimation. These models often incorporate into aging effects and degradation patterns, thereby providing more accurate forecasts of battery performance over their life-time battery_degradation ,Edorado . The temperature is a critical parameter for maintaining battery safety and performance. High temperatures accelerate aging; on the other hand low temperatures can reduce the battery efficiency. The BMS typically includes temperature sensors located at multiple points in the battery pack to monitor temperature fluctuations. Advanced methods for temperature management often deal with combining sensor data with predictive models to optimize the battery’s thermal management system in real-time.
Despite significant progress, several challenges remain in the accurate estimation of SoC and SoH. For example, temperature variations across cells and the aging effects of battery components can complicate estimation processes. Latest innovations in machine learning and adaptive algorithms show promise in improving the accuracy of these estimations machine_SoC . By improving the methods for estimating these parameters, modern BMS contribute significantly to the safety, efficiency, and durability of EV batteries, thereby ensuring that they operate within optimal conditions throughout their lifecycle.
@@ 9 Other Topics and New Trends
The specific energy and power are important parameters in the cell and pack levels that can be countered as performance indices. As mentioned in the previous section, the first step is to construct battery modules as independent units and then integrate them into a pack to fulfill all requirements and standard regulations. The general trend is towards more compact and lighter systems with reduced costs. For pouch cells, an energy density of 650 Wh/L and specific energy of 270 Wh/kg can be found in vehicles. For upcoming cells, a 10-20% enhancement will appear in a couple of years. As an approximation, these numbers can be reduced by approximately 40% at the pack level. For example, an automotive battery pack using cells with 270 Wh/kg can have an energy density of approximately 180 Wh/kg. New cell technologies with improved energy density or specific energy numbers, more integration in packaging, and improved technology for other parts of the system, such as contactors, sensors, and elimination of wires (using wireless), are needed to reduce the weight and volume of the pack. Hence, new materials and chemistry play key roles in meaningful changes in technological advancement. The battery system voltage is another interesting topic. A nominal voltage of 350 V is widely used for passenger cars. However, in some cases, the system voltage was 700-800 V. The power requirement has increased from less than a hundred kW to more than 200 kW, and in some cases, it has increased even further. Consequently, it is necessary to increase the system voltage level to reduce the current level in the system. Consequently, several high-voltage developments, such as 800 V, are under development by the automotive industry, which will increasingly appear in the market Haghbin2 . Another point that should be mentioned is the charging infrastructure.
@@
Another interesting area of research and development in the battery sector involves sodium-based cells, which present an alternative to lithium-ion batteries. Sodium-based batteries utilize sodium ions instead of lithium ions, offering potential advantages such as abundant raw material availability and lower costs, reducing the environmental and ethical concerns associated with lithium and cobalt mining. These sodium-ion batteries are receiving attention for their potential to power EVs in the future, especially as the demand for sustainable and cost-effective solutions. At the begining compared to Li-ion cells, sodium-based cells are expected to play a role in developing batteries which have application in EVs TNO . Another cutting-edge improvements is the development of batteries that do not contain rare earth elements. Many current EV battery chemistries, such as those using LCO battery, depend strongly on rare earth materials. Notwithstanding, as the industry moves toward more sustainable practices, researchers are focusing on new technologies that eliminates or at least minimize the use of these rare earths. These technologies aim to reduce environmental impact and alleviate concerns about supply chain limitations and geopolitical issues surrounding the extraction of these critical materials. TNO battery technology is characterized by high charging and discharging power capabilities. This innovation provides rapid charge/discharge cycles, which offer the potential for reducing charging times significantly and improving vehicle performance. As the automotive industry continues to focus on reducing charging times and enhancing power delivery, TNO’s battery technology could become a key player in meeting these demands TNO_2 . Additionally, battery swapping systems, introduced by NIO, offer a unique solution to the issue of long charging times NIO . Instead of waiting for a long time to charge battery, NIO offers the ability to swap out the battery entirely in a matter of minutes. This system addresses the challenge of limited fast-charging infrastructure and provides a flexible solution for EV owners, particularly in regions with growing EV adoption but underdeveloped charging networks. Increasing the battery voltage may add issues to the available fast-charging infrastructure. Some fast chargers can not charge 800 V batteries. Hence, there are some works under development to change the classical structure such that the vehicle can utilize the low-voltage charger to charge high-voltage packs. In addition to passenger cars, battery packs are required for buses, trucks, and aviation vehicles. Usually, the solutions developed for passenger cars are optimized, and the results are expected to be utilized in other sections. For aviation electrification, some major hindrances require special attention and further development. The classical structure of a battery pack is inappropriate for aviation. It is not possible to disconnect the battery pack in the case of an extreme fault, and new architectures are required. In addition, the environmental conditions are harsher and packaging may be more challenging. Available aviation solutions, such as satellites, are extremely expensive and not affordable. Hence, more work and maturity are required in this sector.
10 Conclusion